OpenStudy (anonymous):

What is the solution of 10 = |7-3x|?

OpenStudy (anonymous):

@Michele_Laino

jimthompson5910 (jim_thompson5910):

`10 = |7-3x|` breaks down into these two equations `10 = 7-3x` or `-10 = 7 - 3x`

jimthompson5910 (jim_thompson5910):

I'm using the rule if |x| = k, then x = k or x = -k where k is some positive number

OpenStudy (anonymous):

x=-1,17/3

OpenStudy (anonymous):

@jim_thompson5910 SO what is it simplified down

jimthompson5910 (jim_thompson5910):

are you able to solve 10 = 7-3x ?